# RobMOT: Robust 3D Multi-Object Tracking by Observational Noise and State Estimation Drift Mitigation on LiDAR PointCloud

### v3

@@ Abstract
This work addresses limitations in recent 3D tracking-by-detection methods, focusing on identifying legitimate trajectories and addressing state estimation drift in Kalman filters. Current methods rely heavily on threshold-based filtering of false positive detections using detection scores to prevent ghost trajectories. However, this approach is inadequate for distant and partially occluded objects, where detection scores tend to drop, potentially leading to false positives exceeding the threshold. Additionally, the literature generally treats detections as precise localizations of objects. Our research reveals that noise in detections impacts localization information, causing trajectory drift for occluded objects and hindering recovery. To this end, we propose a novel online track validity mechanism that temporally distinguishes between legitimate and ghost tracks, along with a multi-stage observational gating process for incoming observations. This mechanism significantly improves tracking performance, with a $6.28\%$ in HOTA and a $17.87\%$ increase in MOTA. We also introduce a refinement to the Kalman filter that enhances noise mitigation in trajectory drift, leading to more robust state estimation for occluded objects. Our framework, RobMOT, outperforms state-of-the-art methods, including deep learning approaches, across various detectors, achieving up to a $4\%$ margin in HOTA and $6\%$ in MOTA. RobMOT excels under challenging conditions, such as prolonged occlusions and tracking distant objects, with up to a 59% improvement in processing latency.
@@
Figure 1: A trajectory drift scenario involving a car (ID4) hidden by a van (ID2). (a) At frame 0, four cars are detected and identified (ID1, ID2, ID3, ID4). (b) By frame 34, the method CastTrack [1] incorrectly estimates the position of car ID4 due to state estimation drift. A dashed green rectangle indicates the actual position. (c) Our proposed approach, with its noise mitigation techniques, successfully localizes car ID4 accurately.
@@
3D multi-object tracking (MOT) is one of the pillars of autonomous driving. It is responsible for determining the state of objects in the environment, including their spatial location and motion parameters like velocity and acceleration. Trajectory drift has been observed in state estimation of occluded objects in the top-performing state-of-the-art CasTrack [1], as shown in Figure 1, failing to recover objects under challenging tracking conditions like prolonged occlusion and distant objects as emphasized in Figure 7. Our investigation shows that this drift results from fluctuation noise attached to the incoming 3D detections from deep learning models, which tracking-by-detection methods rely on for object localization on LiDAR point clouds. This noise impacts the precision of state estimation for spatial location and motion parameters of occluded objects that ultimately leads to deviation from the actual motion trajectory.
@@
Another issue is the absence of trajectory legitimacy verification in the MOT literature [1, 2, 3, 4, 5, 6]. These works apply a fine-tuned threshold on detection scores of the detections provided by deep learning models to filter out false positive observations that cause ghost trajectories. This approach is not reliable as it can block observations that belong to legitimate trajectories while leaking false positive observations. Since observations of legitimate trajectories can inherit some of the false positive characteristics, such as low detection scores for distant and partial occlusion scenarios, the MOT methods need to validate trajectories based on the consistency and legitimacy of their historical observations instead.
@@
In this work, we tackle the above-mentioned challenges through Kalman filter refinement to mitigate the trajectory drift noise, Section 2.2, and propose a novel online mechanism to validate and distinguish between legitimate and ghost trajectories based on studying the characteristics of ghost tracks, Section 2.3. Our proposed refinement shows consistency in tracking for occluded objects superior to the latest state-of-the-art approach, CasTrack, demonstrated in Figure 1 and Figure 7. Next, the proposed trajectory validity mechanism incorporates a multi-stage observational gate that allows potential detections of legitimate tracks to pass to the system and an uncertainty formulation used to quantify the temporal improvement of trajectory certainty based on ghost track characteristics. The mechanism shows a $4\%$ and $6\%$ enhancement in HOTA and MOTA metrics, respectively.
We present the above innovative aspects within a 3D MOT tracking framework dubbed RobMOT. Our contribution can be summarized as follows:
@@
    To the best of our knowledge, this is the first work to uncover the trajectory drift noise associated with detections and its impact on the state estimation of occluded objects in 3D MOT. Our refinement in Kalman filter shows superior performance over the top-performing benchmark, CastTrack [1], on challenging MOT conditions, demonstrated in Figure 7.
@@
    We propose a novel online trajectory validity mechanism for temporal distinguishing between ghost and legitimate tracks, supported by a multi-stage observational gate for potential observation identification of the legitimate tracks. The mechanism reduces about 80% of ghost tracks and enhances HOTA by 7%, as demonstrated in the ablation study Section 3.2.2.
@@
    We propose an online 3D MOT framework with exceptional tracking performance, surpassing recent benchmarks to show its robustness across different datasets and detectors. The tracking performance margin of our framework with the state-of-the-art CasTrack [1] reaches $4\%$ in HOTA and $6\%$ in MOTA, and the enhancement in computational latency is up to $53\%$, as demonstrated in Figure LABEL:fig:hota_mota_comp.
@@ 1 Related Work
Figure 2: The RobMOT framework consists of four modules: a multi-stage observational gate for identifying legitimate observations, an association trajectory innovation for trajectories’ association with observations and next-state prediction, a trajectory validity for temporal validation of the trajectories, and a cache for storing and termination.
@@
3D MOT methods on LiDAR point cloud could be categorized into learning and non-learning (classical) approaches. Part of learning-based methods in the literature involves end-to-end deep learning models that combine object detection and tracking [7, 8, 9, 10]. Meanwhile, other methods adapt tracking-by-detection paradigm [11, 12, 13, 14] when the detection of objects is given and the aim is to perform temporal tracking of the objects. The recent research integrates attention mechanism and transformers concepts [15, 16] into MOT to enhance the tracking performance. However, learning-based solutions require evolved hardware to be applicable to real-time applications.
@@
Meanwhile, classical MOT solutions [4, 5, 2, 3, 6, 17, 1] have gained much attention due to their low latency and the comparative performance with learning-based solutions. Despite recent advancements in the classical methods, challenges such as the limitations of the Kalman filter, ghost tracks, and memory management persist, especially in autonomous driving applications where the environment constantly changes. We will shed light on these challenges and discuss the remedies proposed in the literature next.
@@
Kalman filter (KF) is widely used in tracking-by-detection methods [2, 3, 17, 1] for autonomous driving because of its performance in filtering noise from state estimation while dealing with noisy measurements. If the noise is not properly modeled in the filter, it can cause deviation in the state estimation, as the filter’s estimation is sensitive to imprecise measurements. Cao et al. [6] prevent error accumulation in state prediction caused by periodic occlusion of objects by refining the object’s trajectory after object reappearance. However, the object should be recovered in order to refine its trajectory. Moreover, as of the point cloud’s sparsity, the incoming detections from 3D detectors suffer from spatial distortion around the object’s actual location. This spatial distortion results in an imprecise estimation of the motion parameters, ultimately leading to a deviation in the state estimation once the object faces an occlusion.
@@
Ghost tracks are false positive detections generating illusion tracks that could eventually lead to wrong navigation of the AV. Since the tracking-by-detection paradigm relies heavily on detection accuracy, ghost tracks in object tracking can significantly harm the tracking performance. Kim et al. [3] introduce a multi-stage data association method to overcome this issue by employing multiple detection sources. Nevertheless, they still depend on predefined thresholds to prevent false positives. The downside of a fine-tuned threshold while filtering ghost tracks is the potential of leaking false positive detections that exceed the threshold while blocking detections that belong to actual tracks. Other approaches [5, 2] incorporate a filtration step to address false positive detections. Wang.X et al. [2] confirm the legitimacy of a track when three consecutive detections of that track have been observed. In contrast, Wang.L et al.  [5] select objects for the highest 30% detections in the confidence score. However, the assumptions provided by these solutions may not apply to all scenarios, particularly distant objects or objects subject to periodic occlusions.
Memory management plays a critical role in addressing false positive detections. Maintaining ghost tracks in memory can lead to invalid associations. Meanwhile, early elimination may lose legitimate tracks under prolonged occlusion. In the literature, some work [2, 3, 5, 6] set up a threshold to discard tracks from memory based on how long these tracks have not been observed. Although short memory is suitable for ghost track elimination, it could also remove tracks of objects facing prolonged occlusion. On the other hand, long memory will not remove ghost tracks while increasing the computational time. Nagy et al. [17] attempt to address the trade-off by proposing an MOT solution with low latency and leveraging speed to extend the lifetime of objects, thus capturing more occlusion scenarios. While this approach excels in tracking objects with prolonged occlusion, it does not handle false positives. Wu et al. [1] resolve the issue by associating a confidence score to pre-observed objects in object association. This score decreases over time when the object becomes not observed for a period of time to prevent potential associations with false positive detections. Nonetheless, this approach cannot recover objects under prolonged occlusion as their association score will decrease.
@@
This work proposes RobMOT, a framework to mitigate state estimation deviation caused by fluctuation noise. It also provides adaptive identification for ghost tracks with a quick elimination from memory alongside a mechanism that prevents legitimate track detections from being blocked while filtering ghost tracks.
@@ 2 Methodology
As shown in Figure 2, the framework consists of four parts. A multi-stage observational gate: responsible for identifying potential observations of legitimate tracks. Refined Kalman filter for noise mitigation: a refined KF that updates the state estimation of cached trajectories and predicts their next state. Trajectory validity: Updates the legitimacy uncertainty of trajectories recently associated to observations. Cache: Holds trajectories stored in the framework and discards terminated tracks.
@@ 2.1 Multi-Stage Observational Gate
The multi-stage observational gate is an identification step for potential observations of the legitimate tracks identified by the trajectory validity stage, (see Section 2.3). The primary purpose of this stage is to allow observations of distant and partially occluded objects with low detection scores to enter while preventing overwhelming the system.
There are three types of the observations: observations of existing legitimate tracks, observations of tracks not generated yet by the system, and false observations that cause ghost tracks. Despite ghost tracks identification in the trajectory validity stage, discussed in Section 2.3, a threshold $\alpha_{nconf}$ should be assigned to the incoming observation to prevent overloading the MOT method because 3D deep learning detectors on LiDAR point cloud deliver immense number of detections with various detection scores. Thus, $\alpha_{nconf}$ is assigned a minimal value for detection scores to avoid overloading the system, regardless of the observations’ legitimacy, as this will be assessed later during the trajectory validity stage.
Even though assigning $\alpha_{nconf}$ with a small value allows a high volume of observations to pass to the system, it could block legitimate observations. Hence, We initially identify potential observations of existing legitimate tracks using the same Ecludiean distance threshold $\sigma$ used in the association step (see Section 2.2). Observations located in a distance less than $\sigma$  to a legitimate track, identified by the trajectory validity stage in Section 2.3, are considered legitimate observations. Although the identified potential legitimate observations can pass directly to the system, we have observed that it still causes high latency because of flooding detections from deep learning models on the point cloud. Thus, we provide an additional detection $\alpha_{conf}$ score filtration when $\alpha_{conf}\leq\alpha_{nconf}$ to allow legitimate observations to enter the system while preventing overloading the system. $\alpha_{nconf}$ and $\alpha_{conf}$ are fine-tuned so that no enhancement in the tracking performance is observed since their preliminary purpose is to prevent overloading the system. Lastly, the filtered observations are projected from the LiDAR coordinate to the world coordinate. Algorithm 1 demonstrates these procedures in detail.
@@
     $confirmed\leftarrow false$
@@
           $confirmed\leftarrow true$
@@
2.2 Refined Kalman filter for Noise Mitigation

### v4

@@ Abstract
This paper addresses key limitations in recent 3D tracking-by-detection methods, focusing on the challenges of identifying legitimate trajectories and mitigating state estimation drift in the Kalman filter. Current methods rely heavily on threshold-based filtering of detection scores approach to reduce false positives and prevent ghost trajectories. However, this approach fails for distant and partially occluded objects, where detection scores drop, and false positives surpass that threshold. Additionally, many existing methods assume detections offer precise localization, overlooking inherent noise, which affects localization accuracy and causes state drift for occluded objects, as demonstrated in this work. To this end, a novel track validity mechanism, combined with a multi-stage observational gating process, is proposed that significantly reduces ghost tracks and improves tracking performance. Our method achieves $29.47\%$ enhancement in Multi-Object tracking accuracy (MOTA) on the KITTI validation dataset with the Second detector. Furthermore, a refined Kalman filter term mitigates localization noise, ensuring robust state estimation for occluded objects and superior recovery during prolonged occlusions, with higher order tracking accuracy (HOTA) improving by $4.8\%$ on the KITTI validation dataset with PV-RCNN detector. The proposed online framework, RobMOT, outperforms state-of-the-art methods, including deep learning approaches, across multiple detectors, with HOTA improvements of up to $3.92\%$ on the KITTI testing dataset and $8.7\%$ on the KITTI validation dataset while achieving the lowest identity switch (IDSW) scores of $7$ and 00, respectively. RobMOT excels under challenging scenarios, such as tracking distant objects and handling prolonged occlusions, surpassing state-of-the-art methods on the Waymo Open testing dataset with a $1.77\%$ improvement in MOTA for objects at distances exceeding $50$ meters. RobMOT achieves a groundbreaking run-time of 3221 FPS using a single CPU, establishing itself as a highly efficient and scalable solution for real-time multi-object tracking.
1 Introduction
@@
Figure 1: A trajectory drift scenario involving a car (ID4) hidden by a van (ID2). (a) At frame 0, four cars are detected and identified (ID1, ID2, ID3, ID4). (b) By frame 34, the state-of-the-art method [1, 2] incorrectly estimates the position of car ID4 due to state estimation drift. A dashed green rectangle indicates the actual position. (c) Our proposed approach, with its noise mitigation techniques, successfully localizes car ID4 accurately.
@@
3D multi-object tracking (MOT) is one of the pillars of autonomous driving. It is responsible for determining the state of objects in the environment, including their spatial location and motion parameters like velocity and acceleration. Wu et al. [3, 1], propose an online MOT method, CasTrack, with CasA detector [4] that outperforms the recently published benchmarks, according to the KITTI leaderboard [5]. In another experiment, the author employs the same tracking method with the VirConv detector [2] named VirConvTrack [3]. However, trajectory drift has been observed in state estimation of occluded objects in their work, as shown in Figure 1, preventing MOT methods from recovering objects under challenging tracking conditions like prolonged occlusion and distant objects as emphasized in Figure 8. Our investigation shows that the drift is caused by cumulative inaccuracies in bounding box localization from deep learning models. These models produce detection bounding boxes that slightly shift around an object’s true position in LiDAR data over time. This spatial localization noise creates an illusion of motion, ultimately affecting the object’s state estimation accuracy.
@@
Another issue is the absence of trajectory legitimacy verification in the MOT literature [1, 6, 7, 8, 9, 10]. These works apply a fine-tuned threshold on detection scores of the detections provided by deep learning models to filter out false positive observations that cause ghost trajectories. This approach is not reliable as it can block observations that belong to actual (legitimate) trajectories while leaking false positive observations. Since observations of legitimate trajectories can inherit some of the false positive characteristics, such as low detection scores for distant and partial occlusion scenarios, the MOT methods need to validate trajectories based on the consistency and legitimacy of their historical observations instead.
@@
This work tackles the above-mentioned challenges. First, a refined Kalman filter (KF) is introduced that mitigates the spatial localization noise responsible for trajectory drift in Section 3.2. Second, a novel online mechanism is introduced that validates and distinguishes between legitimate and ghost trajectories based on studying the characteristics of ghost tracks, discussed in Section 3.3. Our proposed refinement shows consistency in tracking for occluded objects superior to the latest state-of-the-art approach [1, 3] demonstrated in Figure 1 and Figure 8. Next, the proposed trajectory validity mechanism incorporates a multi-stage observational gate that allows potential detections of legitimate tracks to pass to the system and an uncertainty formulation used to quantify the temporal improvement of trajectory certainty based on ghost track characteristics. On the KITTI validation dataset [5], the mechanism shows a $8.93\%$ and $29.47\%$ enhancement in higher order tracking accuracy (HOTA) and Multi-Object tracking accuracy (MOTA) metrics with Second detector [11], respectively.
RobMOT is an online 3D MOT framework that integrates the above-mentioned innovations. Our contribution can be summarized as follows:
@@
    To the best of our knowledge, this is the first work to uncover the trajectory drift noise associated with detections and its impact on the state estimation of occluded objects in 3D MOT. Our refinement in KF shows superior performance over the top-performing benchmark [1, 3] on challenging MOT conditions, demonstrated in Figure 8.
@@
    A novel online trajectory validity mechanism is proposed for temporal distinguishing between ghost and legitimate tracks, supported by a multi-stage observational gate for potential observation identification of the legitimate tracks. The mechanism reduces ghost tracks that improve HOTA by $8.93\%$ and MOTA by $29.47\%$ with Second detector [11], as demonstrated in the ablation study Section 4.3.2.
@@
    Lastly, Robmot is an online 3D MOT framework with exceptional tracking performance, surpassing recent benchmarks to show its robustness across different datasets and detectors. The tracking performance margin of our framework with the state-of-the-art [1, 3] reaches $8.7\%$ in HOTA and $16.1\%$ in MOTA on KITTI validation dataset, and significant enhancement in computational latency as demonstrated in Figure 9.
@@ 2 Related Work
Figure 2: An overview diagram of the proposed method (RobMOT) that consists of five modules. A) Multi-stage observational gate that prevents detections from overloading the framework while allowing detections that potentially belong to confirmed trajectories to pass. B) Trajectory association that associates trajectories stored in the cache with current detections. C) State estimation is used to update and predict the following spatial state of trajectories after the association stage while mitigating the spatial localization noise. D) Trajectory validity mechanism that updates the temporal validity score for trajectories associated with new observations. E) Cache that maintains active trajectories while terminating trajectories with high uncertainty about their estimated spatial state.
@@
3D MOT methods on LiDAR point can be categorized into learning and non-learning (classical) approaches. Part of learning-based methods in the literature involves end-to-end deep learning models that combine object detection and tracking [12, 13, 14, 15, 16]. Meanwhile, other methods adapt tracking-by-detection paradigm [17, 18, 19, 20] when the detection of objects is given and the aim is to perform temporal tracking of the objects. The recent research integrates attention mechanism and transformers concepts [21, 22] into MOT to enhance the tracking performance. However, learning-based solutions require evolved hardware to be applicable to real-time applications.
@@
Meanwhile, classical MOT solutions [8, 9, 6, 7, 10, 23, 1, 24, 25] have gained much attention due to their low latency and the comparative performance with learning-based solutions. Despite recent advancements in the classical methods, challenges such as the limitations of the KF, ghost tracks, and memory management remain, particularly in autonomous driving applications where the environment constantly changes. This section will shed light on these challenges and discuss solutions adopted in the literature.
@@
Kalman filter is widely used in tracking-by-detection methods [6, 7, 23, 1] for autonomous driving because of its performance in filtering noise from state estimation while dealing with noisy measurements. If the noise is not properly modeled in the filter, it can cause deviation in the state estimation, as the filter’s estimation is sensitive to imprecise measurements. Cao et al.[10] prevent error accumulation in state prediction caused by periodic occlusion of objects by refining the object’s trajectory after object reappearance. However, they need to recover the object to refine its trajectory. Recent work in vehicle tracking[26] has shown that integrating probabilistic models and driver preview characteristics can improve state estimation and prediction, offering potential insights for enhancing Kalman filter-based tracking, especially in scenarios involving occlusion. Due to the point cloud’s sparsity, our research shows that detections from deep learning detectors on the 3D LiDAR point cloud suffer from spatial distortion around the actual location of the objects (Section 3.2). This spatial distortion results in drift in state estimation for occluded objects as their motion parameters estimation becomes imprecise.
@@
Ghost tracks are false positive detections generating illusion tracks that could eventually lead to wrong navigation of the autonomous vehicle (AV). Since the tracking-by-detection paradigm relies heavily on detection accuracy, ghost tracks in object tracking can significantly harm the tracking performance. Kim et al. [7] introduce a multi-stage data association method to overcome this issue by employing multiple detection sources. Similarly, Zhang et al. [25] employs a two-stage association by prioritizing the association of detections with a high confidence score in the first stage and the second stage for a low confidence score to prevent potential ghost tracks from being associated with valid tracks. Nevertheless, they still depend on predefined thresholds to prevent false positives. The downside of a fine-tuned threshold while filtering ghost tracks is the potential of leaking false positive detections that exceed the threshold while blocking detections that belong to actual tracks. Other approaches [9, 6, 24] incorporate a filtration step to address false positive detections. Wang.X et al. [6] confirm the legitimacy of a track when three consecutive detections of that track have been observed. In contrast, Wang.L et al.  [9] select objects for the highest 30% detections in the confidence score. However, the assumptions provided by these solutions may not apply to all scenarios, particularly distant objects or objects subject to periodic occlusions.
Stadler and Beyerer [24], on the other hand, attempt to identify ghost tracks and filter them out by excluding bounding boxes that have high intersection over union (IoU) with others; however, this approach has two limitations; the first drawback is that it can exclude non-ghost track detections in congestion scenarios while it assumes ghost tracks are always generated from duplicated detections for the same object even though Figure 10 shows a ghost track (Bounding box with ID:50) that appears far from other objects, as a false detection from the detector.
Memory management plays a critical role in addressing false positive detections and object occlusion. It is most common in MOT methods [6, 7, 23, 1] to extend the lifetime of tracks in the memory to handle occlusions, particularly prolonged ones. However, this expansion comes with the cost of late elimination for ghost tracks that harm the tracking performance.
Wu et al. [1] resolve the issue by associating a confidence score to pre-observed objects in object association. This score decreases over time when the object becomes not observed for a period of time to prevent potential associations with false positive detections. Nonetheless, this approach cannot recover objects under prolonged occlusion as their association score will decrease.
@@
This work proposes RobMOT, a framework to mitigate state estimation deviation caused by the spatial localization noise from deep learning detectors. It also provides an online adaptive identification mechanism for ghost tracks integrated with a multi-stage observational gate that prevents legitimate track detections from being blocked. The framework handles the trade-off between fast removal for ghost tracks and recovering prolonged occlusions using uncertainty-based memory, which allows quick elimination for ghost tracks while extending the lifetime of legitimate tracks.
@@ 3 Methodology
As shown in Figure 2, the framework comprises five modules. A) Multi-stage observational gate: prevents overwhelming the framework with unnecessary observations while allowing potential valid observations to pass. B) Trajectory association: associates trajectories in the Cache with current detections. C) State estimation: a refined KF that updates state estimation of trajectories and predicts their next state while mitigating the spatial localization noise of the employed detector. D)Trajectory validity mechanism: update the temporal validity score for trajectories with new observations. E) Cache: holds active trajectories in the system while terminating those with high uncertainty about their spatial state estimation.
@@ 3.1 Multi-Stage Observational Gate
Although the trajectory validity mechanism stage (Section 3.3) is responsible for discrimination between ghost and legitimate tracks, detectors on the 3D point cloud provide an enormous number of detections that can significantly impact the run time of the framework. Hence, the multi-stage observational gate is introduced to limit the overflow of the detections and allow detection that potentially belongs to legitimate tracks to pass even with a low detection score.
The incoming detections can be categorized into three types: observations that belong to legitimate tracks in the Cache, new tracks, and false observations that lead to the formation of ghost tracks. Initially, a threshold $\alpha_{nconf}$ is assigned to prevent overloading the MOT method by detections; however, this threshold is assigned a minimal detection score value that prevents the detections overflow, regardless of false observations, since ghost tracks are handled later by the trajectory validity mechanism.
Although assigning $\alpha_{nconf}$ with a small value allows a high volume of observations to pass through the system, it can block observations belonging to legitimate tracks in the memory. Hence, observations that potentially belong to legitimate tracks in the cache are identified using the same Euclidean distance threshold $\sigma$ used in the association step (see Section 3.2).Observations with a distance less than $\sigma$  to a legitimate track are identified as potential legitimate observations.
Although potential legitimate observations can pass directly to the system, they still cause high latency because of flooding detections from deep learning models on the point cloud. Thus, an additional detection $\alpha_{conf}$ score filtration is provided such that $\alpha_{conf}\leq\alpha_{nconf}$ to allow legitimate observations to enter the system while preventing overloading the system. $\alpha_{nconf}$ and $\alpha_{conf}$ are fine-tuned so that no enhancement in the tracking performance is observed since their preliminary purpose is to prevent overloading the system. Lastly, the filtered observations are projected from the LiDAR coordinate to the world coordinate.
To summarize, the multi-stage observational gate identifies potential legitimate detections, donated by $n$, from the utilized detector using legitimate trajectories stored in the memory, donated by $l$, that contains $m$ trajectories such that $l\leq m$. In the worst case, this module’s run-time complexity is $O(nl)$. Algorithm 1 demonstrates these procedures in detail.
@@
     $det.confirmed\leftarrow false$
     if $det$.score $<\alpha_{nconf}$ then
@@
              $det.confirmed\leftarrow true$
@@
3.2 Trajectory Association and State Estimation with a Refined Kalman Filter for Noise Mitigation
In the trajectory association stage, the permitted observations are associated with trajectories cached in the framework. The Hungarian algorithm [27] is used with the algebraic association strategy followed in [23]. The association for an observation is performed by selecting a trajectory with the shortest Euclidean distance between the estimated state and the observation. When the shortest distance to an observation exceeds $\sigma$, a new trajectory is created for the observation.
In the next stage (State estimation), the framework performs next-state estimation for all trajectories cached in the system. Trajectories associated with an observation will pass through state update and state prediction as demonstrated in Figure 2. Other trajectories will go directly through state prediction. The framework utilizes KF with a constant acceleration motion model so that the state space of a trajectory consists of its spatial location, velocity, and acceleration in $x$ and $y$ directions.